OpenStudy (anonymous):

HELP!!! I WILL GIVE MEDALS!!! :) <3 What decimal number is equal to 7/9?

OpenStudy (anonymous):

do they have answer choices?

OpenStudy (anonymous):

Oh yes..sorry.. Let me type them.

OpenStudy (anonymous):

There are a few ways to go about this kind of problem. Remember that decimals are just fractions over 10 or 100 or 1,000, etc. So you can make the fraction have a denominator (the bottom part of the fraction) become one of those numbers. For example 3/5 5 * 2 = 10 and 3 * 2 = 6 So, 3/5 is the same as 6/10 or six-tenths, which is written 0.6 in decimal form. Another way is to divide: For example 1 / 3. Well, 3 doesn't go into 1, but it does go into 10, so we instead could write 1.0 / 3 3 goes into 10 3 times with 1 left over. But we are actually using 1.0 instead of 10 so we get 0.3 with .1 left over. If we try dividing 3 into 100 instead we would get 100 / 3 which gives us 33 and one left over But since that was 1.00 we actually get 0.33 with .01 left over. If you have a calculator, you can do this division in the calculator, and you will get 0.3333333 (with 3's all the way across the screen). This division method can be used for all the fractions. using a calculator (or doing it by hand) you will find 1/3 = 0.33333333 1/6 = 0.16666666 7/8 = 0.875 3/5 = 0.6 5/8 = 0.625 2/3 = 0.66666666 7/9 = 0.77777777 6/12 = 0.5 hope that helps

OpenStudy (perl):

http://www.wolframalpha.com/input/?i=7%2F9

OpenStudy (anonymous):

0.7

OpenStudy (anonymous):

Okay..thanks...

OpenStudy (anonymous):

isnt it .7?

OpenStudy (anonymous):

I already failed my test, oh well.

OpenStudy (anonymous):

-.- Im sorry :'(

OpenStudy (anonymous):

...

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

but did I get your question right if I did can I get my medal

OpenStudy (anonymous):

THANKS !!! sorry for your test :(

OpenStudy (anonymous):

Get out of here you stupid Jason.